# Machine-learned flow estimation with sparse data — exemplified for the rooftop of a UAV vertiport

### v1

@@ Abstract
We propose a physics-informed data-driven framework for urban wind estimation. This framework validates and incorporates the Reynolds number independence for turbulent flows, thus allowing the extrapolation for wind conditions far beyond the training data. Another key enabler is a machine-learned non-dimensionalized manifold from snapshot data. The velocity field is modeled using a double encoder-decoder approach. The first encoder normalizes data using the oncoming wind speed, while the second encoder projects this normalized data onto the isometric feature mapping manifold. The decoders reverse this process, with $k$-nearest neighbor performing the first decoding and the second undoing the normalization. The manifold is coarse-grained by clustering to reduce the computational load for de- and encoding. The sensor-based flow estimation is based on the estimate of the oncoming wind speed and a mapping from sensor signal to the manifold latent variables. The proposed machine-learned flow estimation framework is exemplified for the flow above an Unmanned Aerial Vehicle (UAV) vertiport. The wind estimation is shown to generalize well for rare wind conditions, not included in the original database.
@@ I Introduction
Weerasuriya et al. (2018) Tsuen Wan in Hong Kong EXP (Wind tunnel)
Ricci et al. (2017) Livorno in Italy EXP (Wind tunnel)
Gao et al. (2012) Building complex EXP (On-site sensor)
Zou et al. (2021) Darlington campus of University of Sydney EXP (On-site sensor)
@@
Research on urban wind estimation primarily falls into three categories: computational fluid dynamics (CFD) simulations, experiments (wind tunnel testing, field measurements), and data-centric methods(Gao et al., 2023). Table 1 summarizes representative research. CFD methods can provide full-state information under varying wind conditions, such as wind angle and velocity. For example, Toparlar et al. (2015) conducted unsteady Reynolds-averaged Navier–Stokes (uRANS) simulations on wind flow and temperatures in the Bergpolder Zuid region, demonstrating CFD’s potential for accurately predicting urban microclimates. Similarly, Ramponi et al. (2015) used 3D steady RANS simulations on generic urban configurations, highlighting the effects of wind angle and urban morphology on wind field patterns. Jiang et al. (2024) introduced a methodology for defining no-fly zones (NFZs) for drones using CFD, offering three hazardous indices—safe, deviation, and unsafe—to assess drone operation conditions. Additionally, Van Hooff and Blocken (2010) used CFD to evaluate the influence of wind angle and urban geometry on air exchange rates within a stadium model. Simplified generic geometries, often composed of regular arrays of obstacles, are also commonly used for systematic studies of aerodynamic and geometric correlations, as they are computationally cheaper (Buccolieri, Sandberg, and Di Sabatino, 2010; Hang and Li, 2010; Moonen, Dorer, and Carmeliet, 2011; Lin et al., 2014). Wind tunnel experiments and on-site measurements likewise provide valuable insights into urban wind patterns, with notable methods including wind tunnel and fan-array wind generator experiments(Weerasuriya et al., 2018; Ricci et al., 2017; Wang et al., 2024; Li et al., 2024; Liu et al., 2024), as well as the on-site wind data collected by sensors(Gao et al., 2012; Zou et al., 2021).
@@
Figure 2: An example of flow estimation with extrapolation, exemplified for the rooftop of UAV vertiport. (a) Testing data (red dot), the nondimensionalized database (black dots) and the nearest neighbor of the testing data (red star) displayed in a 2D manifold. (b) Estimated flow field. (c) Velocity field of the nearest neighbor. (d) Velocity field of the testing data. For testing sensor signals outside the existing database, the nearest neighbor may yield inaccurate flow information. Instead, we use the estimated oncoming wind speed to non-dimensionalize the testing data and subsequently restore the estimated flow field. By incorporating Reynolds number independence, this approach effectively generalizes to rare wind conditions.
@@ II.1 Numerical simulation
The fluid flow is simulated by solving the non-dimensionalized incompressible Reynolds-averaged Navier-Stokes (RANS) equations. The characteristic scales used for non-dimensionalization are the height of the vertiport $H=2.8$ m, the free stream velocity $U_{\infty}$ and twice the dynamic pressure $\rho U_{\infty}^{2}$. The velocity field is noted $\bm{u}=(u,v,w)$, $p$ is the static pressure, $Re=U_{\infty}H/\nu$ is the Reynolds number and $\nu$ is the kinematic viscosity. One RANS simulation is performed for each case, i.e., for each wind condition. A commercially available CFD software, ANSYS Fluent 15.0, is employed to discretize the governing equations via the cell-centered Finite Volume Method (FVM). Spatial discretization of the governing equations is conducted using a second-order scheme, while the temporal term is discretized using a first-order implicit scheme. For the boundary conditions, a uniform streamwise velocity is imposed at the inlet, defined as $\bm{u}=(U_{\infty},0,0)$. At the outlet, an outflow condition is applied, which imposes a Neumann condition on the velocity, $\partial_{x}\bm{u}=(0,0,0)$, and a Dirichlet condition on the pressure, $p=0$. A no-slip boundary condition is enforced on the surface of the vertiport. Additionally, interface boundary conditions are applied to the connecting regions between the inner and outer domains. Slip boundary conditions are used on the outer domain, ensuring that any wake-wall interactions are excluded.
To create a CFD database as training data, the wind conditions include nine levels based on the Beaufort Scale $b$, with wind angles varying from $0^{\circ}$ to $360^{\circ}$. Specifically, the wind speeds are denoted as $b_{0}=0.2$ m/s, $b_{1}=1.5$ m/s, $b_{2}=3.3$ m/s, $b_{3}=5.4$ m/s, $b_{4}=7.9$ m/s, $b_{5}=10.7$ m/s, $b_{6}=13.8$ m/s, $b_{7}=17.1$ m/s, and $b_{8}=20.7$ m/s, with wind angles incremented by $3^{\circ}$, resulting in a total of $1080$ simulation cases.
II.2 Validation
@@
Grids Inner domain Outer domain Number of points
1 362 mio 101 mio 463 mio
2 543 mio 101 mio 644 mio
3 543 mio 195 mio 738 mio
4 543 mio 282 mio 835 mio
5 734 mio 282 mio 1016 mio
@@ III Machine-learned flow estimation with sparse data
Figure 6: Methodology framework. In the modeling phase, the dataset is non-dimensionalized based on the oncoming wind speed. The non-dimensionalized velocity fields serve as training inputs for the manifold model, using the ISOMAP-$k$NN manifold learner. In the estimation phase, a wind condition estimator first predicts the oncoming wind speed when a testing sensor signal is available. The sensor signals are non-dimensionalized based on this estimated speed and input into the manifold learner. After mapping these signals to latent variables, the decoder reconstructs the complete velocity fields.
@@ III.1 Non-dimensionalization of the data
Similarly, for the velocity sensor data $s$, the non-dimensionalization is:
@@ III.2 Clustering
Since the initial cluster centroids are randomly generated, the $k$-means++ algorithm is repeated 100 times in this study to ensure optimal partitioning. The averaged centroids $\bm{c}_{k}$ and the cluster affiliation $k({m})$ of the snapshots are taken as the final results.
@@ III.3 Non-dimensionalized manifold learner
We develop a non-dimensionalized manifold learner based on the approach of  Farzamnik et al. (2023), as shown in figure 7. The proposed manifold learner consists of two main parts. First, in the modeling part, the high-dimensional data is embedded into a low-dimensional space using isometric feature mapping (ISOMAP). This data-driven encoding process uncovers a hidden manifold, allowing us to connect the new coordinates to physical features of the dataset, such as wind angle and wind speed around the vertiport. Subsequently, a decoding process is included in the estimation part, enabling the latent variables to return to the high-dimensional space and reconstruct the original velocity fields.
@@ III.3.1 Encoder
The $k$-nearest neighbor ($k$NN) regression algorithm (Stone, 1977; Fix, 1985) is the foundation for constructing the manifold model. The $k$NN algorithm is a non-parametric approach to approximate a mapping function $f$ by averaging the outputs of the $k$-nearest observations to the input data, as shown in figure 8. Choosing a smaller $k$ can make the algorithm sensitive to outliers and noise, while a larger $k$ generally increases the computational cost. For a velocity field $\bm{u}^{m}$, we define $\mathcal{N}(\bm{u}^{m};K,\bm{u}^{1},\ldots,\bm{u}^{M})$ as the $K$ nearest neighbors of this velocity field in the dataset $\bm{u}^{1},\ldots,\bm{u}^{M}$.
@@ III.4 Flow estimation from sensor signal
$\hat{U}_{\infty}(t)=\lambda^{l}(t)U_{\infty}^{l}.$ (24)
@@
$\sum_{m=1}^{M}\chi^{*}_{t}(m)=K_{w}.$ (30)
Where $\mathcal{N}(\bm{s}^{*}(t);K_{w},\bm{s}^{*1},\ldots,\bm{s}^{*M})$ is a set of $K_{w}$ nearest neighbors to $\bm{s}^{*}(t)$ in $\bm{s}^{*1},\ldots,\bm{s}^{*M}$, $\bm{w}^{*}_{m}(t)$ is subsequently defined as:
@@ IV Results
Figure 10: (a) The discretized manifold from the centroids and (b) the projected full manifold from the snapshots.
@@
Figure 13: Manifolds of the velocity sensor signal. The same $2$-D manifold as in the flow field data suggests that the flow characteristics are captured. The overall shape appears slightly square, which is consistent with the rectangular distribution of the sensor locations around the vertiport.
@@ IV.1 Flow manifold on the rooftop of a UAV vertiport
Figure 10 illustrates the phase portrait of the coarse-grained centroid manifold and the projected full manifold. Here, $\gamma_{1}$ and $\gamma_{2}$ represent the $2$ latent variable cooridinates. It is worth noting that the values of $\gamma$ are determined by the norm between datasets, which varies based on configurations and has no specific physical interpretations. Thus, the focus should be on their relative values and relationships between different variables. Both manifolds exhibit a circular structure, indicating symmetry and clear periodicity within the dataset. This circular shape suggests that the underlying dynamics of the system are regular and predictable, consistent with the periodic wind angle information contained in the data. The snapshot manifold preserves the general circular shape of the centroid manifold while incorporating finer details, resulting in a smoother manifold shape. These additional details may capture subtle variations within the data, offering a refined view of the system’s behavior. However, the manifold is not entirely uniform or continuous. This irregularity stems from the geometric configuration of the vertiport, which leads to a non-uniform distribution of centroids that affects the shape and continuity of the full snapshot manifold. Consequently, certain areas may appear more compressed or stretched.
@@
To facilitate subsequent flow estimation, the sensor signals are also taken for modeling. This step helps identify data patterns from the sensors and verify their validity for flow estimation, though it is not essential to the overall framework. Due to the smaller size of the sensor dataset, the snapshot manifold can be directly obtained with $K_{e}=20$. Figure 12 displays the residual variance of the velocity sensor, revealing a similar $2$-D manifold as seen in the velocity field data. This similarity indicates that the flow characteristics are accurately captured, confirming the effectiveness of the current sensor placement and the comprehensiveness of the sensor signals. As shown in the sensor signal manifolds in Figure 13, the shape remains circular and symmetric. The overall shape appears slightly square, reflecting the rectangular distribution of the sensor locations around the vertiport. Additionally, the circular shape of the manifold allows for the extrapolation of higher wind speeds through non-dimensionalization, enabling the estimation of rare wind conditions. This capability to predict under unusual scenarios, such as wind gusts, highlights the robustness of the manifold-based approach.
@@ IV.2 Flow estimation on the rooftop of a UAV vertiport from sensor signals
Table 4: The estimation error of the oncoming wind speed and velocity field by different cases from the velocity sensor. For details, see text.
Case Oncoming velocity Manifold type Data type Estimation type Error $\hat{U}_{\infty}$ (%) Error (%)

### v2

# Machine-learned flow estimation with sparse data — exemplified for the rooftop of an unmanned aerial vehicle vertiport

@@ Abstract
We propose a physics-informed data-driven framework for urban wind estimation. This framework validates and incorporates the Reynolds number independence for flows under various working conditions, thus allowing the extrapolation for wind conditions far beyond the training data. Another key enabler is a machine-learned non-dimensionalized manifold from snapshot data. The velocity field is modeled using a double encoder-decoder approach. The first encoder normalizes data using the oncoming wind speed, while the second encoder projects this normalized data onto the isometric feature mapping manifold. The decoders reverse this process, with $k$-nearest neighbor performing the first decoding and the second undoing the normalization. The manifold is coarse-grained by clustering to reduce the computational load for de- and encoding. The sensor-based flow estimation is based on the estimate of the oncoming wind speed and a mapping from sensor signal to the manifold latent variables. The proposed machine-learned flow estimation framework is exemplified for the flow above an Unmanned Aerial Vehicle vertiport. The wind estimation is shown to generalize well for rare wind conditions, not included in the original database.
@@ I Introduction
Weerasuriya et al. (2018) Tsuen Wan in Hong Kong Experimental (Wind tunnel)
Ricci et al. (2017) Livorno in Italy Experimental (Wind tunnel)
Gao et al. (2012) Building complex Experimental (On-site sensor)
Zou et al. (2021) Darlington campus of University of Sydney Experimental (On-site sensor)
@@
Research on urban wind estimation primarily falls into three categories: computational fluid dynamics (CFD) simulations, experiments (wind tunnel testing, field measurements), and data-centric methods(Gao et al., 2023). Table 1 summarizes representative research. CFD methods can provide full-state information under varying wind conditions, such as wind angle and velocity. For example, Toparlar et al. (2015) conducted unsteady Reynolds-averaged Navier–Stokes (uRANS) simulations on wind flow and temperatures in the Bergpolder Zuid region, demonstrating CFD’s potential for accurately predicting urban microclimates. Similarly, Ramponi et al. (2015) used 3D steady RANS simulations on generic urban configurations, highlighting the effects of wind angle and urban morphology on wind field patterns. Jiang et al. (2024) introduced a methodology for defining no-fly zones for drones using CFD, offering three hazardous indices—safe, deviation, and unsafe—to assess drone operation conditions. Additionally, Van Hooff and Blocken (2010) used CFD to evaluate the influence of wind angle and urban geometry on air exchange rates within a stadium model. Simplified generic geometries, often composed of regular arrays of obstacles, are also commonly used for systematic studies of aerodynamic and geometric correlations, as they are computationally cheaper (Buccolieri, Sandberg, and Di Sabatino, 2010; Hang and Li, 2010; Moonen, Dorer, and Carmeliet, 2011; Lin et al., 2014). Wind tunnel experiments and on-site measurements likewise provide valuable insights into urban wind patterns, with notable methods including wind tunnel and fan-array wind generator experiments(Weerasuriya et al., 2018; Ricci et al., 2017; Wang et al., 2024; Li et al., 2024; Liu et al., 2024), as well as the on-site wind data collected by sensors(Gao et al., 2012; Zou et al., 2021).
@@
Figure 2: An example of flow estimation with extrapolation, exemplified for the rooftop of UAV vertiport. (a) Testing data (red dot), the nondimensionalized database (black dots) and the nearest neighbor of the testing data (red star) displayed in a 2D representation of the manifold. (b) Estimated flow field. (c) Velocity field of the nearest neighbor. (d) Velocity field of the testing data. For testing sensor signals outside the existing database, the nearest neighbor may yield inaccurate flow information. Instead, we use the estimated oncoming wind speed to non-dimensionalize the testing data and subsequently restore the estimated flow field. By incorporating Reynolds number independence, this approach effectively generalizes to rare wind conditions.
@@ II.1 Numerical simulation
The fluid flow is simulated by solving the non-dimensionalized incompressible Reynolds-averaged Navier-Stokes equations. The characteristic scales used for non-dimensionalization are the height of the vertiport $H=2.8$ m, the free stream velocity $U_{\infty}$ and twice the dynamic pressure $\rho U_{\infty}^{2}$. The velocity field is noted $\bm{u}=(u,v,w)$, $p$ is the static pressure, $Re=U_{\infty}H/\nu$ is the Reynolds number and $\nu$ is the kinematic viscosity. One RANS simulation is performed for each case, i.e., for each wind condition. A commercially available CFD software, ANSYS Fluent 15.0, is employed to discretize the governing equations via the cell-centered Finite Volume Method. Spatial discretization of the governing equations is conducted using a second-order scheme, while the temporal term is discretized using a first-order implicit scheme. For the boundary conditions, a uniform streamwise velocity is imposed at the inlet, defined as $\bm{u}=(U_{\infty},0,0)$. At the outlet, an outflow condition is applied, which imposes a Neumann condition on the velocity, $\partial_{x}\bm{u}=(0,0,0)$, and a Dirichlet condition on the pressure, $p=0$. A no-slip boundary condition is enforced on the surface of the vertiport. Additionally, interface boundary conditions are applied to the connecting regions between the inner and outer domains. Slip boundary conditions are used on the outer domain, ensuring that any wake-wall interactions are excluded.
To create a CFD database as training data, the wind conditions include nine levels based on the Beaufort scale $b$, with wind angles varying from $0^{\circ}$ to $360^{\circ}$. Specifically, the wind speeds are denoted as $b_{0}=0.2$ m/s, $b_{1}=1.5$ m/s, $b_{2}=3.3$ m/s, $b_{3}=5.4$ m/s, $b_{4}=7.9$ m/s, $b_{5}=10.7$ m/s, $b_{6}=13.8$ m/s, $b_{7}=17.1$ m/s, and $b_{8}=20.7$ m/s, with wind angles incremented by $3^{\circ}$, resulting in a total of $1080$ simulation cases. Four testing datasets are created to evaluate the performance of the model. In the first dataset, a single wind speed of $U_{\infty}=b_{4}=7.9$ m/s is chosen, with all associated wind angles from the CFD database, resulting in a total of $120$ snapshots. The second dataset consists of randomly selected snapshots within the CFD database, with multiple wind speeds chosen from $b_{0}=0.2$ m/s to $b_{8}=20.7$ m/s and wind angles from $0^{\circ}$ to $360^{\circ}$, totaling $216$ snapshots. Additional simulations have been carried out for the third dataset. The working conditions are randomly chosen within the database range, i.e., wind speeds are randomly distributed between $0.2$ and $20.7$ m/s and wind angles between $0^{\circ}$ and $360^{\circ}$, totaling $216$ snapshots. The fourth dataset contains new simulations with working conditions extending beyond the database range. Wind speeds are randomly distributed between $0.2$ and $30$ m/s, and wind angles are also randomly distributed between $0^{\circ}$ and $360^{\circ}$, totaling $216$ snapshots.
II.2 Verification
@@
Grids Inner domain (million) Outer domain (million) Number of points (million)
1 362 101 463
2 543 101 644
3 543 195 738
4 543 282 835
5 734 282 1016
@@ III Machine-learned flow estimation with sparse data
Figure 6: Methodology framework. Modeling: The dataset is first non-dimensionalized based on the oncoming wind speed. The non-dimensionalized velocity fields serve as training inputs for the manifold model, using the ISOMAP-$k$NN manifold learner. Estimation: The wind condition estimator first predicts the oncoming wind speed for a given testing sensor signal with index $t$. The sensor signals are then non-dimensionalized based on this estimated speed and input into the manifold learner. After mapping these signals to latent variables, the manifold model reconstructs the complete velocity fields.
@@ III.1 Non-dimensionalization of the data
Similarly, for the velocity sensor data $\bm{s}$, the non-dimensionalization is:
@@ III.2 Clustering
Based on Li et al. (2021b), we set the parameters for the $k$-means++ algorithm to $1000$ iterations and $100$ repetitions to ensure optimal partitioning. The averaged centroids $\bm{c}_{k}$ and the cluster affiliation $k({m})$ of the snapshots are taken as the final results.
@@ III.3 Non-dimensionalized manifold learner
We develop a non-dimensionalized manifold learner based on the approach of  Farzamnik et al. (2023), as shown in figure 7. The proposed manifold learner consists of two main parts. First, in the modeling part, the high-dimensional data is embedded into a low-dimensional space using isometric feature mapping. This data-driven encoding process uncovers a hidden manifold, allowing us to connect the new coordinates to physical features of the dataset, such as wind angle and wind speed around the vertiport. Subsequently, a decoding process is included in the estimation part, enabling the latent variables to return to the high-dimensional space and reconstruct the original velocity fields.
@@ III.3.1 Encoder
The $k$-nearest neighbor regression algorithm (Stone, 1977; Fix, 1985) is the foundation for constructing the manifold model. The $k$NN algorithm is a non-parametric approach to approximate a mapping function $f$ by averaging the outputs of the $k$-nearest observations to the input data, as shown in figure 8. Choosing a smaller $k$ can make the algorithm sensitive to outliers and noise, while a larger $k$ generally increases the computational cost. For a velocity field $\bm{u}^{m}$, we define $\mathcal{N}(\bm{u}^{m};K,\bm{u}^{1},\ldots,\bm{u}^{M})$ as the $K$ nearest neighbors of this velocity field in the dataset $\bm{u}^{1},\ldots,\bm{u}^{M}$.
@@ III.4 Flow estimation from sensor signal
$\hat{U}_{\infty}(t)=\lambda^{l}(t)U_{\infty}^{l},$ (24)
@@
$\sum_{m=1}^{M}\chi^{*}_{t}(m)=K_{w},$ (30)
Where $\mathcal{N}(\bm{s}^{*}(t);K_{w},\bm{s}^{*1},\ldots,\bm{s}^{*M})$ is the set of $K_{w}$ nearest neighbors to $\bm{s}^{*}(t)$ in $\bm{s}^{*1},\ldots,\bm{s}^{*M}$. $\bm{w}^{*}_{m}(t)$ is subsequently defined as:
@@ IV Results
Figure 10: (a) The centroid-based manifold and (b) the snapshot-based manifold, with the representative flow state of $U_{\infty}=20.7$ m/s at $z=0.2$ m visualized around the manifold. $\gamma_{1}$ and $\gamma_{2}$ represent the latent variable coordinates. There is a slight shift in the current feature coordinates with respect to the wind angle, causing the distribution to be asymmetric. This shift is determined by the MDS process in the encoder of the manifold leaner, which focuses on the manifold shape and may generate the latent variables on rotated feature coordinates if the manifold is axisymmetric.
@@ IV.1 Flow manifold on the rooftop of a UAV vertiport
Figure 10 illustrates the phase portrait of the coarse-grained centroid manifold and the projected full manifold. Here, $\gamma_{1}$ and $\gamma_{2}$ represent the $2$ latent variable coordinates. It is worth noting that the values of $\gamma$ are determined by the norm between datasets, which varies based on configurations and has no specific physical interpretations. Thus, the focus should be on their relative values and relationships between different variables. Both manifolds exhibit a circular structure, indicating symmetry and clear periodicity within the dataset. This circular shape suggests that the underlying dynamics of the system are regular and predictable, consistent with the periodic wind angle information contained in the data. The snapshot manifold preserves the general circular shape of the centroid manifold while incorporating finer details, resulting in a smoother manifold shape. These additional details may capture subtle variations within the data, offering a refined view of the system’s behavior. However, the manifold is not entirely uniform or continuous. This irregularity stems from the geometric configuration of the vertiport, which leads to a non-uniform distribution of centroids that affects the shape and continuity of the full snapshot manifold. Consequently, certain areas may appear more compressed or stretched.
@@
Figure 13: Manifold of the velocity sensor signals, with the oncoming wind angles of the large dots plotted around the manifold. The manifold shape is similar to the snapshot data manifold, suggesting that the flow characteristics are captured.
The sensor signals are also taken for modeling. This step helps identify data patterns from the sensors and verify their validity for flow estimation, though it is not essential to the overall framework. Due to the smaller size of the sensor dataset, the snapshot manifold can be directly obtained with $K_{e}=20$. Figure 12 displays the residual variance of the velocity sensor, revealing a similar $2$-D manifold as seen in the velocity field data. This similarity indicates that the flow characteristics are accurately captured, confirming the effectiveness of the current sensor placement and the comprehensiveness of the sensor signals. As shown in the sensor signal manifolds in Figure 13, the shape remains circular and symmetric. The overall shape appears slightly square, reflecting the rectangular distribution of the sensor locations around the vertiport. Additionally, the circular shape of the manifold allows for the extrapolation of higher wind speeds through non-dimensionalization, enabling the estimation of rare wind conditions. This capability to predict under unusual scenarios, such as wind gusts, highlights the robustness of the manifold-based approach.
@@ IV.2 Flow estimation on the rooftop of a UAV vertiport from sensor signals
Figure 14: Estimation error of test dataset 4: The estimation error of the oncoming velocity $\hat{U}_{\infty}$ is shown in (a) for wind speed $U_{\infty}$ and in (b) for wind angle $\alpha$. The normalized squared error (representation error) of the velocity field is presented in (c) against wind speed $U_{\infty}$ and in (d) against wind angle $\alpha$. The normalized mean squared error is highlighted in red ($E=6.55\%$), while the blue dot indicates the snapshot for further visualization.
Figure 15: Estimated velocity field using the velocity sensor. The top left shows the slice locations. The dashed line in the flow visualizations indicates the position of the vertiport. The arrow starting from the bottom left point on the visualizations indicates the wind angle. Top: CFD dataset. Middle: estimated field. Bottom: the absolute estimation error. Snapshot wind conditions: $U_{\infty}=26.55$ m/s, $\alpha=65.22^{\circ}$, with an estimation error of $E=2.25\%$.
Figure 14 shows the estimation error of $\hat{U}_{\infty}$ and $\hat{u}$ for all test wind conditions in testing dataset 4 (data for extrapolation). The error distribution is relatively uniform across various wind speeds and angles, indicating no specific conditions with significantly higher errors, which demonstrates the generalization capability of the framework. The most significant error among all estimated velocity fields is within $10\%$, while most have errors below $5\%$. Even for the extrapolated wind conditions which are not included in the CFD database, the error distribution remains consistent with the rest of the dataset. An estimated snapshot, marked in blue and close to the normalized mean squared error, is selected for flow visualization in Figure 15. The absolute estimation error $\left|\bm{u}-\hat{\bm{u}}\right|$ is plotted in the bottom row. This visualization includes the estimated flow field across four rooftop transverse planes at heights of $z=0.2$ m, $0.4$ m, $0.6$ m, and $0.8$ m. The accurate prediction of the wind angle is evident, with the flow structure in the estimated data closely resembling that of the CFD data, highlighting the framework’s effectiveness in capturing full-state flow dynamics.
IV.3 Error sources of the model
Table 4: Estimation error of the oncoming wind speed and velocity field by different cases from the velocity sensor. For details, see text.
Case Oncoming velocity Manifold type Data type Estimation type $\hat{U}_{\infty}$ Error (%) Error (%)